A -3 microcoulomb charge is at the origin and a +4 microcoulomb charge is at a distance of x=2m away. Find the point between the charges where the voltage is 0.
Answer #1

let the point be x ,

then , as due to a point charge , potential is given as

V = KQ/d

Here, at the point x

K*(-3/x + 4/(2-x)) = 0

then

4x = 6 - 3x

x = 6/7

x = 0.857 m

the point where V = 0 is x = 0.857 m
Answer #2


electric potential due to a point charge is given by V = KQ/r


wjhere Q is charge and Ris distance

so here

let V be zero at xunits fron Q1


so at this point

V1 = V2


KQ/(x)   = Kq2/(2-x)

2-x/x   = 4/3   = 1.333

2-x = 1.33x

x = 2/2.66

x = 0.857 m or 85.7 cm   to the right of Q1 net V is zero